# The impact of data set similarity and diversity on transfer learning success in time series forecasting

### v1

@@ Abstract
Models, pre-trained on a similar or diverse source data set, have become pivotal in enhancing the efficiency and accuracy of time series forecasting on target data sets by leveraging transfer learning. While benchmarks validate the performance of model generalization on various target data sets, there is no structured research providing similarity and diversity measures explaining which characteristics of source and target data lead to transfer learning success. Our study pioneers in systematically evaluating the impact of source-target similarity and source diversity on zero-shot and fine-tuned forecasting outcomes in terms of accuracy, bias, and uncertainty estimation. We investigate these dynamics using pre-trained neural networks across five public source datasets, applied in forecasting five target data sets, including real-world wholesales data. We identify two feature-based similarity and diversity measures showing: Source-target similarity enhances forecasting accuracy and reduces bias, while source diversity enhances forecasting accuracy and uncertainty estimation and increases the bias.
@@
Transfer learning, Deep neural networks, Foundation models
@@ 1 Introduction
Transfer learning for time series forecasting describes pre-training a model on one or more source data sets in order to either continue training on the target data set (fine-tuning) or directly produce forecasts (zero-shot) (Pan and Yang,, 2009). Studies of the last few years have shown that transfer learning can successfully enhance prediction performance or reduce computational and data requirements, compared to training a model from scratch on the target data. Research can be divided into two approaches: pre-train on a similar source data set from the application perspective (specifically pre-trained model) to forecast a specific target data set, e.g. Oreshkin et al., (2021)Wellens et al., (2022) or pre-train on a huge and diverse set of time series (foundation model (Bommasani et al.,, 2022)) to supply a general foundation for forecasting any target data set, e.g. Das et al., (2024)Goswami et al., (2024).
@@
While the attributes of source-target similarity and source diversity drive the research on transfer learning for time series forecasting, existing contributions explain little on how the choice of the source data set impacts transfer learning success, i.e. better zero-shot or fine-tuned forecasting performance than from scratch, on diverse target data sets. Especially if source and target data sets stem from different domains, transfer learning success might look like a lucky strike, leaving it unclear whether it can be transmitted to other use cases. But still, there are little studies comparing the similarity of different source data sets to the targets. One of them is He et al., (2019) who show that source data sets which are similar in Dynamic Time Warping (DTW) to the target, provide better fine-tuned accuracy. However, feature-based similarity measures, used to compare sources of different domains in regard to their target forecasting performance, are under-researched. While foundation models are built on massive diverse source data sets, there is no standardized definition of data set diversity and papers comparing sources of different diversity for transfer learning are lacking.
@@
In contrast, most studies which analyze transfer learning results, concentrate on model architectures, hyper-parameters and training procedures and this only either for the zero-shot or fine-tuned case, e.g. Oreshkin et al., (2021)Kambale et al., (2023)Feng et al., (2024). A possible reason for this lack might be the seemingly limited choice of potential similar source data sets in applications and the method of selecting as many source data sets as possible to built a foundation data set. However, it is important to fill these gaps to provide more explainability to pre-trained models and formulate indications for the success of transfer learning in new use cases.
@@
To achieve this, we first pre-train the probabilistic recurrent neural network DeepAR (Salinas et al.,, 2020) on five publicly available source data sets of different domains and sizes. Additionally, we pre-train it on a concatenation of all these sets, comprising around 7.5 million data points, to create a small foundation model with respect to the training set, which we call the Multisource model. We apply the pre-trained models in a zero-shot and fine-tuned manner to forecast five target data sets, of which three are public and two are sales data, originating from different wholesalers. We measure the resulting performance in terms of forecasting accuracy, bias, uncertainty estimation and training time and benchmark the performance to the statistical model ETS. To identify data set characteristics, we analyze the data at hand by calculating two different feature sets. Based on these as well as using a shape-based approach, we quantify similarities between the sources and targets. The features also serve as basis for the computation of source data diversity, which we define as the variance of time series features of the series within a data set. Additionally we display data set diversity as time series features in the latent PCA space. Eventually, we map the forecasting performance on the target data to the similarity and diversity results: the higher the source diversity or source-target similarity, the more accurate the forecasts on the target. For the bias, source-target similarity leads to its reduction, while the source diversity increases it. Prediction intervals are more accurate, if the source data is more diverse. These relationships are much stronger for the zero-shot than the fine-tuned forecasting results. We did not identify any consistent relation between forecasting performance and the regarded shape-based similarity and diversity metrics.
@@
    we pioneer in providing a structured investigation on how the source-target similarity and source diversity relates to zero-shot and fine-tuned transfer learning forecasting accuracy, bias and uncertainty estimation for data sets of different domains.
@@
    we provide the code used in this study open source (upon acceptance), containing methods for transfer learning, feature and shape-based measures for calculating data set characteristics, relations between source-target similarity and source diversity and transfer learning performance as well as the pre-trained models to enable reproduction of the results and encourage further research.
The rest of the paper is structured as follows. First, we look at related work from the two lines of research, specific pre-trained and foundation models, highlighting successful transfer learning and their source and target data set motivations. In Section 3 we describe the experimental design of our transfer learning study, including models, data sets and metrics. After that, we show the results in terms of accuracy, bias, uncertainty estimation and training times in Section 4. To examine data set characteristics, we present similarity and diversity measures and their concrete values for the data at hand in Section 5. The relationship between these measures and forecasting performance is evaluated in Section 6. We summarize and show directions for further research in the conclusion in Section 7.
@@ 2 Related work
While conversational AI models are the result of significant research on foundation models for natural language processing, the discussion about transfer learning and pre-trained models frequently omits time series analysis. Only in recent years, the interest in transfer learning has conveyed to time series – for a review we refer to Weber et al., (2021). In the following, we summarize related work for time series analysis and forecasting, showing the potential of pre-trained models and the motivation behind the selected source and target data sets. While some literature focuses on finding a similar source data set for a given target data set, another stream aims to provide a general foundation model.
@@ 2.1 Specifically pre-trained models
Pre-trained models have particular justification in forecasting applications, when the training data set is small, for example, when forecasting demand for a product that has just been introduced to the market. Karb et al., (2020) show that a model pre-trained for a similar product can improve accuracy and enhance uncertainty estimates compared to the scratch model trained on a new product of a food retailer. Miraftabzadeh et al., (2023) forecast the 24 hours energy production of a new PV plant with a fine-tuned LSTM pre-trained on a nearby plant’s data, which works better than training from scratch or zero-shot forecasting.
@@
Apart from the reduced data requirements, pre-trained models can also save training time. Wellens et al., (2022) show that the M5 competition winning LightGBM architecture pre-trained on more aggregated hierarchies and only fine-tuned at less aggregated ones of the M5 data set, can save 25% of training time while keeping similar accuracy levels, compared to training from scratch.
Furthermore, pre-trained models can also be competitive in forecasting accuracy of normal sized data sets. Oreshkin et al., (2021) demonstrate that a pre-trained N-BEATS (Oreshkin et al.,, 2020) on the M4 data or FRED data is competitive against statistical methods for forecasting various data sets without even fine-tuning. They train one source model for each granularity of the M4 data and mostly choose the source model according to the target granularity.
@@
Given a specific target data set, modelers aiming for pre-training seek appropriate, similar source data sets. In that, similarity is often measured regarding apparently common features, e.g., the application domain, the data set, or the granularity. Less papers examine similarities in the shape or the statistical characteristics of the time series within the source and the target data sets. In time series classification, pre-trained models on data sets of similarly shaped time series within the UCR archive (Chen et al.,, 2015), measured by the DTW, help target classification accuracy (Ismail Fawaz et al.,, 2018). For small target data sets, transferring knowledge of data within the same domain help the classification task, even if the pre-training task is different (Ismail-Fawaz et al.,, 2024)He et al., (2019) show an analogical result for financial forecasting: source data sets consisting of two similar data sets, according to the DTW distances to the target data, provide more accurate fine-tuned forecasts than those combined from dissimilar sources. Furthermore, sources from the same industry have a positive effect on the accuracy, as well as using two instead of only one source data sets.
@@
Pre-training with similar source data is also possible with more than one data set. This can be done by either combining the data to one big source and pre-training the model on it (we call it Multisource model) or by pre-training one model per data set and combining the models to an ensemble (we call it Ensemble model). Gu and Dai, (2021) for example, weight the pre-trained models per source according to the target similarity measured by the Maximum Mean Discrepancy and inter source data dependencies to gain a weighted ensemble.
The transfer learning experiment of Woo et al., (2023) with source and target data sets of the same data generating process, shows increased forecasting accuracy and uncertainty estimation on the target for source data sets with more time series, while longer time series lead to similar samples and hence to over-fitting on the source. Therefore, source data diversity plays a crucial role in the success of transfer learning, which we discuss in the following.
@@ 2.2 Foundation models
A general source data set, containing huge amounts of diverse time series, cannot provide as much similarity to a specific target data set as a selected source data set can. But, due to the so called double descent, which describes the test accuracy first improving, then decreasing and again improving with growing model complexity, training data and runtime, a large foundation model can make sense (Nakkiran et al.,, 2021). It brings the advantage of making the search for a specific source data set redundant, with the goal of providing accurate and fast forecasts for any target data. However, there has long been no foundation model for forecasting, which might have been due to the lack of a massive structured public time series repository, like those existing for text and images, to train a large source model (Hahn et al.,, 2023).
@@
The first foundation model for time series forecasting, which gained a lot of attention, was the proprietary TimeGPT-1 (Garza and Mergenthaler-Canseco,, 2023), a transformer based deep learning architecture, pre-trained on a vast amount of time series data from diverse domains. TimeGPT-1 reveals superior performance on monthly and weekly data and also good performance on daily and hourly data. Yet, the authors do not reveal their training and test data sets and stay reserved about the details of their architecture and training parameters. More transparent studies have emerged on foundation models trained on the Monash Time Series Repository (Godahewa et al.,, 2021), which encompasses real-world time series data of diverse domains. The Lag-Llama transformer model of Rasul et al., (2024) e.g., which was pre-trained on it, beats all the scratch benchmark models and another pre-trained transformer, when zero-shot forecasting the held out Traffic (Lai et al.,, 2018) data set. The authors show that the zero-shot performance of Lag-Lama on the held out data sets increases with the number of samples used for pre-training. Additionally, they visualize the diversity of the source data sets by plotting the mean catch22 (Lubba et al.,, 2019) feature vector per data set in the PCA space. Data sets of the same domain are projected into similar regions, suggesting that source data sets consisting of different domains, are more diverse. They call for further research on how the source diversity relates to the forecasting performance, which we adress in this paper. Ekambaram et al., (2024) also used the Monash repository to pre-train the first publicly available forecasting foundation model, which can be used for general time series forecasting. Specifically, it allows for multivariate forecasting, the integration of external variables and different granularities. In their benchmark, the advantage of the foundation model compared to the scratch model decreases with growing target training data and training time.
@@
Foundation models can also be built out of general and diverse synthetically generated time series. In one example,  Dooley et al., (2024) generate 300,000 time series containing multiplicative trend, seasonal and error components to train a transformer model for zero-shot forecasting. The resulting model outperforms the benchmark models like naive, transformer-based scratch models as well as the pre-trained NBEATS model, in terms of the MSE wins over data sets and forecast horizons on seven benchmark data sets. In time series classification, Rotem et al., (2022) created 15 mio time series on which they pre-trained multiple regression task. After fine-tuning for classification on target data from the UCR archive, this source model works better than selecting a source data set from the UCR archive, when regarding only 10% for training or fine-tuning on the target.
@@
Overall, research suggests that the success of a foundation model depends source wise on the size and the diversity of the source data. Target data sets benefit from foundation models, especially if there is limited data or training time. The improved performance of the foundation model compared to the scratch model might be due to a smoother training loss curve on the target data (Yeh et al.,, 2023).
@@
This study extends the existing research by structurally examining the relations between source diversity, source-target similarity and forecasting performance on the target, integrating models pre-trained on single source data sets, their ensemble and also a small foundation model trained on the combination of five data sets from diverse domains.
3 A transfer learning experiment
@@
Given the established state of the art, we first present an experiment designed to explore the forecasting performance of pre-trained models, where the source and target data stem from different domains. To that end, we select five diverse source and target data sets and divide them each into training and test set. We use the training set of each source to pre-train source models and additionally train the Multisource Model on the concatenation of the sources’ training sets. Note that this training data set is relatively small in comparison to those of state-of-the-art foundation models, since we aim to analyze the relation of the data set characteristics to the forecasting performance and limit the computational effort. Still, it can be seen as a small foundation model. Also, we build an ensemble of the source models, the Ensemble model. We evaluate the pre-trained models with regard to their forecasting performance on the test sets of the target data sets for the zero-shot and fine-tuned case. For the former, we directly forecast with the pre-trained models and for the latter we first fine-tune the models on the target training sets. We benchmark the forecasting performance of the pre-trained models to the so-called scratch model, which uses the training data of the target data sets to train from scratch, i.e. with a random start initialization. This comparison provides information about the added value of transfer learning. Also, we include the ETS models as benchmark. The details of the experiment are given in the following.

### v2

@@ Abstract
Pre-trained models have become pivotal in enhancing the efficiency and accuracy of time series forecasting on target data sets by leveraging transfer learning. While benchmarks validate the performance of model generalization on various target data sets, there is no structured research providing similarity and diversity measures to explain which characteristics of source and target data lead to transfer learning success. Our study pioneers in systematically evaluating the impact of source-target similarity and source diversity on zero-shot and fine-tuned forecasting outcomes in terms of accuracy, bias, and uncertainty estimation. We investigate these dynamics using pre-trained neural networks across five public source datasets, applied to forecasting five target data sets, including real-world wholesales data. We identify two feature-based similarity and diversity measures, finding that source-target similarity reduces forecasting bias, while source diversity improves forecasting accuracy and uncertainty estimation, but increases the bias.
@@
Transfer learning, Neural networks, Foundation models, Time series, Prediction intervals
@@
[inst1] organization=Fraunhofer IIS, Fraunhofer Institute for Integrated Circuits IIS, Fraunhofer Center for Applied Research on Supply Chain Services SCS, country=Germany
@@
[inst2]organization=Christian-Albrechts-Universität zu Kiel, Research Group Service Analytics, country=Germany
[…]
[inst3]organization=Otto-Friedrich-Universität Bamberg, Supply Chain Management, country=Germany
[…]
[inst4]organization=Vives University of Applied Science, Business Management, country=Belgium
[…]
[inst5] organization=University of Haute-Alsace, IRIMAS, country=France
[…]
[inst6]organization=Monash University, Department of Data Science and Artificial Intelligence, country=Australia
@@ 1 Introduction
Transfer learning for time series forecasting describes pre-training a model on one or more source data sets in order to either continue training on a target data set (fine-tuning) or directly produce forecasts (zero-shot) (Pan and Yang, 2009). Recent studies have shown that transfer learning can successfully enhance prediction performance or reduce computational and data requirements, compared to training a model from scratch on the target data. The related references can be divided into two approaches: pre-train on a similar source data set from the application perspective (specifically pre-trained model) to forecast a specific target data set, e.g. Oreshkin et al. (2021)Wellens et al. (2022), or pre-train on a huge and diverse set of time series (foundation model (Bommasani et al., 2022)) to supply a general foundation for forecasting any target data set, e.g. Das et al. (2024)Goswami et al. (2024).
@@
Source-target similarity and source diversity are selection criteria for the source data sets of pre-trained models in time series forecasting, see for example (Oreshkin et al., 2021) and (Rasul et al., 2024). Though, existing contributions explain little on how the choice of the source data set impacts transfer learning success, i.e., better zero-shot or fine-tuned forecasting performance than from scratch, on diverse target data sets. Especially if source and target data sets stem from different domains, transfer learning success might look like a lucky strike, leaving it unclear whether it can be transmitted to other use cases. Few existing studies compare the similarity of different source data sets to the targets.  He et al. (2019), who show that source data sets that favor the target in Dynamic Time Warping (DTW) provide better fine-tuned accuracy. However, to the best of our knowledge, no published study yet examines the relation between feature-based similarity measures (Alcock et al., 1999) regarding time series characteristics e.g. trend or auto-regression of series in the source and target data sets, and the transfer learning forecasting performance. While foundation models are built on massive diverse source data sets, there is no standardized definition of data set diversity and papers comparing sources of different diversity for transfer learning are lacking.
@@
In contrast, most studies which analyze transfer learning results, focus on model architectures, hyper-parameters and training procedures and this either for the zero-shot or fine-tuned case (e.g. Oreshkin et al. (2021)Kambale et al. (2023)Feng et al. (2024)). A possible reason for this lack might be the seemingly limited choice of potential similar source data sets in applications and the method of selecting as many source data sets as possible to built a foundation data set. However, it is important to fill these gaps to provide more justification to pre-trained models and formulate indications for the success of transfer learning in new use cases. This can then be used to build targeted public time series databases.
Figure 1: Summary of our approach and research question.
@@
Figure 1 summarizes our research question and approach. First, we pre-train a probabilistic recurrent neural network DeepAR (Salinas et al., 2020) on five publicly available source data sets of different domains and sizes. Additionally, we pre-train it on a concatenation of all these sets, comprising around 7.5 million data points, to create a small foundation model with respect to the training set, which we call the Multisource model. We apply the pre-trained models in a zero-shot and fine-tuned manner to forecast five target data sets, of which three are public and two are non-public sales data, originating from different wholesalers. We measure the resulting performance in terms of forecasting accuracy, bias, and uncertainty estimation and benchmark the performance with ETS (Hyndman et al., 2024). To identify data set characteristics, we analyze the data at hand by calculating two different feature sets. Based on these as well as using a shape-based approach, we quantify similarities between the sources and targets. The features also serve as basis for the computation of source data diversity, which we define as the variance of time series features of the series within a data set. Additionally, we display data set diversity as time series features in the latent principal component analysis (PCA) space. Then, we map the forecasting performance on the target data to the similarity and diversity results using linear regression. We find that higher source diversity in catch22 features leads to more accurate point forecasts and uncertainty estimates on the target data set. Source-target similarity in tsfresh features leads to bias reduction, while the source diversity in tsfresh features increases it. These relationships are much stronger for the zero-shot than the fine-tuned forecasting results. We did not identify any consistent relation between forecasting performance and the regarded shape-based similarity metrics.
@@
    we are the first to provide a structured investigation on how the source-target similarity and source diversity relates to zero-shot and fine-tuned transfer learning forecasting accuracy, bias and uncertainty estimation for data sets of different domains.
@@
    we provide the code used in this study open source (upon acceptance), containing methods for transfer learning, feature- and shape-based measures for calculating data set characteristics, relations between source-target similarity and source diversity and transfer learning performance as well as the pre-trained models to enable reproduction of the results and encourage further research.
The rest of the paper is structured as follows. First, we look at related work from the two lines of research, specifically pre-trained and foundation models. We highlight successful transfer learning experiments and the respective motivations behind the source and target data set selection. In Section 3, we describe the experimental design of our transfer learning study, including models, data sets and metrics. After that, we show the results in terms of accuracy, bias, and uncertainty estimation in Section 4. Then we present the similarity and diversity for the data at hand in Section 5, and evaluate their relationship to forecasting performance in Section 6. In the conclusion we summarize and show directions for further research (Section 7).
@@ 2 Related work
While the recent success of conversational AI models is the result of research on foundation models for natural language processing, the discussion about transfer learning and pre-trained models frequently omits time series analysis. Only in recent years, the interest in transfer learning has conveyed to time series – for a review see Weber et al. (2021). In the following, we summarize related work for time series analysis and forecasting, showing the potential of pre-trained models and the motivation behind the selected source and target data sets. While one stream of literature focuses on finding a similar source data set for a given target data set, the other stream aims to provide a general foundation model.
@@ 2.1 Specifically pre-trained models
Pre-trained models have particular justification in forecasting applications, when the training data set is small, for example when new products or technical plants are introduced: In food retails, pre-trained models on a similar product can improve accuracy and enhance uncertainty estimates compared to the scratch model of a new product (Karb et al., 2020). The forecast of the 24 hours energy production of a new photovoltaic plant with a fine-tuned LSTM pre-trained on a nearby plant’s data, is superior to training from scratch or zero-shot forecasting (Miraftabzadeh et al., 2023). Furthermore, pre-trained models can also be competitive in forecasting accuracy of normal sized data sets. Oreshkin et al. (2021) demonstrate that a pre-trained N-BEATS (Oreshkin et al., 2020) on the M4 data or FRED data is competitive against statistical methods for forecasting various data sets without even fine-tuning. They train one source model for each frequency of the M4 data and mostly choose the source model according to the target frequency.
@@
Apart from the reduced data requirements, pre-trained models can also save training time: Wellens et al. (2022) show that the M5 competition winning LightGBM architecture pre-trained on more aggregated hierarchies and only fine-tuned at less aggregated ones of the M5 data set, can save 25% of training time while keeping similar accuracy levels, compared to training from scratch.
@@
Given a specific target data set, modelers aiming for pre-training seek appropriate, similar source data sets. Similarity is thereby often measured by apparently common features, like the application domain e.g. (Karb et al., 2020), the data set e.g. (Wellens et al., 2022), or the frequency e.g. (Oreshkin et al., 2021). Less papers examine similarities in the shape or the statistical characteristics of the time series within the source and the target data sets. In time series classification, pre-trained models on data sets of similarly shaped time series within the UCR archive (Chen et al., 2015), measured by the DTW, help target classification accuracy (Ismail Fawaz et al., 2018). For small target data sets, transferring knowledge of data within the same domain help the classification task, even if the pre-training task is different (Ismail-Fawaz et al., 2024)He et al. (2019) show an analogical result for financial forecasting: source data sets consisting of two similar data sets, according to the DTW distances to the target data, provide more accurate fine-tuned forecasts than those combined from dissimilar sources. Furthermore, sources from the same industry have a positive effect on the accuracy, as well as using two instead of only one source data sets.
@@
Pre-training with similar source data is also possible with more than one data set. This can be done by either combining the data to one big source and pre-training the model on it (we call it Multisource model) or by pre-training one model per data set and combining the models to an ensemble (we call it Ensemble model). Gu and Dai (2021) for example, weight the pre-trained models per source according to the target similarity measured by the Maximum Mean Discrepancy and inter source data dependencies to gain a weighted ensemble.
While pre-trained models can increase target forecasting accuracy and uncertainty estimation, a small source data set can also lead to over-fitting, which decreases zero-shot performance on the target (Woo et al., 2023). In addition to using a source model of different complexity, pre-training on another similar data set of different size (Nakkiran et al., 2021) or fine-tuning foundation models offer a remedy.
@@ 2.2 Foundation models
Foundation models are trained on a huge and diverse set of data with the aim to build a foundation for any downstream task. Accordingly, selecting a specific source data set is unnecessary. However, until recently the success of time series foundation models was limited, which might have been due to the lack of a massive structured public time series repository, like those existing for text and images, to train a large source model (Hahn et al., 2023).
@@
The first foundation model for time series forecasting, which gained a lot of attention, was the proprietary TimeGPT-1 (Garza and Mergenthaler-Canseco, 2023), a transformer based deep learning architecture, pre-trained on a vast amount of time series data from diverse domains. TimeGPT-1 reveals superior performance on monthly and weekly data and good performance on daily and hourly data. Yet, the authors do not reveal their training and test data sets and stay reserved about the details of their architecture and training parameters. More transparent studies have emerged on foundation models trained on the Monash Time Series Repository (Godahewa et al., 2021), which encompasses real-world time series data of diverse domains. The Lag-Llama transformer model of Rasul et al. (2024) e.g., which was pre-trained on it, beats all the statistical and machine learning benchmark models and another pre-trained transformer, when zero-shot forecasting the held-out Traffic (Lai et al., 2018) data set. The authors show that the zero-shot performance of Lag-Llama on the held out data sets increases with the number of samples used for pre-training. Additionally, they visualize the diversity in time series characteristics of the source data sets by plotting the mean catch22 (Lubba et al., 2019) feature vector per data set in the PCA space. Data sets of the same domain are projected into similar regions, suggesting that source data sets consisting of different domains, are more diverse. They call for further research on how the source diversity relates to the forecasting performance, and we address this in our work. Ekambaram et al. (2024) also used the Monash repository to pre-train the first publicly available forecasting foundation model, which can be used for general time series forecasting. Specifically, it allows for multivariate forecasting, the integration of external variables and different frequencies. Their pre-trained zero and few-shot models show comparable or superior performance against state-of-the-art pre-trained large language models, pre-trained forecasting models and traditionally trained machine learning models on uni- and multi-variate data sets with and without exogenous variables which were not part of the pre-training.
@@
Foundation models can also be built out of general and diverse synthetically generated time series. For example Dooley et al. (2024) generate 300,000 time series containing multiplicative trend, seasonal and error components to train a transformer model for zero-shot forecasting. The resulting model outperforms the benchmark models like naive, transformer-based scratch models as well as the pre-trained NBEATS model, in terms of the MSE on different forecast horizons and seven benchmark data sets. Using pre-trained models on artificial data can be better than searching for an existing source data set. This can also be the case if the pre-training task is different to the target task: Rotem et al. (2022) created 15 million time series on which they pre-trained a CNN for time series regression. After fine-tuning for classification on target data from the UCR archive, this source model works better than selecting a source data set from the UCR archive, when regarding only 10% for training or fine-tuning on the target.
@@
Overall, research suggests that the success of a foundation model depends source wise on the size and the diversity of the source data. Foundation models are specifically competitive when there is limited data in the target data set or training time.
@@
This study expands the current research by structurally examining the relationships between source diversity, source-target similarity, and forecasting performance on the target. It integrates models pre-trained on single source data sets, their ensemble, and a small foundation model trained on the combination of five data sets from diverse domains.
3 Transfer learning experiment
@@
We first explore the forecasting performance of pre-trained models, where the source and target data stem from different domains. To that end, we select five diverse source and target data sets and divide them each into training and test set. We use the training set of each source to pre-train source models and additionally train the Multisource model on the concatenation of the sources’ training sets. Note that this training data set is relatively small in comparison to those of state-of-the-art foundation models, since we aim to analyze the relation of the data set characteristics to the forecasting performance and limit the computational effort. Still, it can be seen as a small foundation model. Furthermore, we build an ensemble of the source models, the Ensemble model. We evaluate the pre-trained models with regard to their forecasting performance on the test sets of the target data sets for the zero-shot and fine-tuned case. For the former, we directly forecast with the pre-trained models and for the latter we first fine-tune the models on the target training sets. We benchmark the forecasting performance of the pre-trained models against the so-called Scratch model, which uses the training data of the target data sets to train from scratch, i.e. with a random start initialization. This comparison provides information about the added value of transfer learning. Also, we include the ETS models as benchmark. The details of the experiment are given in the following.